Session: Nov 2002 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Huang Xinyong at Thu Nov 21 15:51:36 2002) Minor failures :
The RCP receiver is broken down. So, we only have the LCP for the experiment. We start it at UT: 10:30:35.

Urumqi (zhanghua at Thu Nov 21 19:03:49 2002) Success : HIGH
It is snow. 16:40:00--16:51:00(UT),the source exceed threshold of antenna.

Medicina (Roda J. at Thu Nov 21 22:07:16 2002) Success :

Jodrell_Bank (Alastair Gunn at Fri Nov 22 01:05:35 2002) Minor failures :
Some source changes had inadequate slew times. Late on source for several scans. No other known problems.

Cambridge (Alastair Gunn at Fri Nov 22 01:06:03 2002) Success :
No known problems.

Onsala (Michael Lindqvist at Fri Nov 22 08:11:37 2002) Success :
No known problems. Onsala only have LCP.

Metsahovi (Ari Mujunen at Mon Nov 25 12:24:11 2002) Minor failures : RECO
Weather is quite bad, it is snowing constantly during first few hours, last 4h or so are reasonable. Please note that we have a new H-maser, new cabling. Don't expect too much from high-numbered tape tracks, high parity errors expected.